OpenStudy (anonymous):

Simpify3^3 √375 15^3 √3 15^3 √15 8^3 √3 5^3 √3

OpenStudy (anonymous):

from the choices listed we can deduce a couple things

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when simplifying a radical it looks like this\[\sqrt{\text{\large number}}=\sqrt{\text{perfect \square }\times \text{imperfect \square}}\]

OpenStudy (anonymous):

if there is \[\sqrt{3} \space \text{Then we know} \sqrt{3\times 125}\]would be the reduced radical. But 125 is not a perfect square and cannot be brought outside.

OpenStudy (anonymous):

\[\sqrt{A}=\sqrt{a^2\times b}=a \sqrt{b}\]

OpenStudy (anonymous):

\[\sqrt{375}=\sqrt{15*25}=\sqrt{15}\sqrt{25}=5\sqrt{15}\]

OpenStudy (anonymous):

would the answer be 5^3√3 ?

OpenStudy (anonymous):

options r wrong :)